# Special pure gradings on simple Lie algebras of types $E_{6}$, $E_{7}$, $E_{8}$

### v1

@@ 4. Special pure gradings on simple Lie algebras of types $E_{6}$, $E_{7}$, or $E_{8}$
Up to equivalence, there are exactly five special pure gradings on a simple Lie algebra of type $E_{8}$ over an algebraically closed field of characteristic 00. Their universal groups and types are given in the next table:
@@
The quadratic form $q\colon\overline{R}\rightarrow\mathbb{F}_{2}$ in (2.2) is regular with trivial Arf invariant. Moreover, the nonisotropic vectors of $\overline{R}$ are precisely the $120$ vectors $\bar{\alpha}:=\alpha+2R$ for $\alpha\in\Phi^{+}$. By Proposition 3.2, the complementary lattice $E$ satisfies that $\overline{E}=E/2R$ is a totally isotropic subspace of $\overline{R}$, and hence it has dimension at most $4$. Moreover, for any dimension up to $4$, any two totally isotropic subspace of $\overline{R}$ are conjugate under the action of $\mathrm{O}(q)$. As a consequence, we get the following possibilities, according to (3.6):
@@
On the other hand, there are $120-24=96$ positive roots $\alpha\in\Phi^{+}$ with $b_{q}(\bar{\alpha},\overline{E})\neq 0$, and then the left coset $\bar{\alpha}+E$ contains just two nonisotropic vectors, thus obtaining $48$ pairs of positive roots, which give $96$ homogeneous components in (3.6) of dimension $2$. Hence, the type is $(0,96,0,12,0,0,0,1)$.
@@
Up to equivalence, there are exactly four special pure gradings on a simple Lie algebra of type $E_{7}$ over an algebraically closed field of characteristic 00. Their universal groups and types are given in the next table:
@@
The quadratic form $q\colon\overline{R}\rightarrow\mathbb{F}_{2}$ in (2.2) is regular, but odd-dimensional. The radical of $b_{q}$ is one-dimensional: $\operatorname{\mathrm{rad}}b_{q}=\mathbb{F}_{2}a$, with $q(a)=1$. Moreover, the nonisotropic vectors of $\overline{R}$ are precisely the $63$ vectors $\bar{\alpha}:=\alpha+2R$ for $\alpha\in\Phi^{+}$ together with $a$. Any subspace of $\overline{R}$ containing strictly $\operatorname{\mathrm{rad}}b_{q}$ contains other nonisotropic vectors besides $a$, so it is not of the form $\overline{E}=E/2R$ with $E$ the complementary lattice of a special pure grading. Also, the lattice $F$ with $\overline{F}=\operatorname{\mathrm{rad}}b_{q}$ cannot be the complementary lattice of a special pure grading, graded by its universal group, because for any two different positive roots $\alpha,\beta\in\Phi^{+}$, $\bar{\alpha}-\bar{\beta}\neq a$, as otherwise we would have $01=q(\bar{\beta})=q(\bar{\alpha}+a)=q(\bar{\alpha})+q(a)=1+1=$ 0, and then (3.4) is not fulfilled. Therefore, the complementary lattice $E$ of a special pure grading satisfies that $\overline{E}$ is a totally isotropic subspace of $\overline{R}$ and, in particular, it does not contain $\operatorname{\mathrm{rad}}b_{q}$.
@@
Here the quadratic space $\overline{E}^{\perp}/\overline{E}$ is regular of dimension $5$, so it contains $16$ nonisotropic elements, one of them being $a+\overline{E}$. For each $\alpha\in\Phi^{+}$ with $\bar{\alpha}\in\overline{E}^{\perp}$, the coset $\bar{\alpha}+\overline{E}$ consists of two nonisotropic vectors, so either $\bar{\alpha}+\overline{E}=\{\bar{\alpha},\bar{\beta}\}$ for another positive root $\beta$, or $\bar{\alpha}+\overline{E}=\{\bar{\alpha},a\}$, which happens only for $\bar{\alpha}=a+u$, with $\overline{E}=\mathbb{F}_{2}u$. Hence there are $15$ pairs of positive roots that come together in (3.6). The other $35=63-30$ positive roots give rise to one-dimensional homogeneous components and the type is $(66,30,0,0,0,0,1)$.
@@
In this case the quadratic space $\overline{E}^{\perp}/\overline{E}$ is regular of dimension $3$, so it contains $4$ nonisotropic elements, one of them being $a+\overline{E}$. For each $\alpha\in\Phi^{+}$ with $\bar{\alpha}\in\overline{E}^{\perp}$, the coset $\bar{\alpha}+\overline{E}$ consists of four nonisotropic vectors, which come from four positive roots unless $\bar{\alpha}+\overline{E}=a+\overline{E}$ that contains three nonisotropic vectors obtained from positive roots. For the remaining $48=63-15$ positive roots $\alpha$ with $\bar{\alpha}\not\in\overline{E}^{\perp}$, the coset $\bar{\alpha}+\overline{E}$ contains two nonisotropic vectors. The type is then $(0,48,2,6,0,0,1)$.
@@
For the remaining $56=63-7$ positive roots $\alpha$ with $\bar{\alpha}\not\in\overline{E}^{\perp}$, the coset $\bar{\alpha}+\overline{E}$ contains four nonisotropic vectors. The type is then $(0,0,0,28,0,0,3)$. ∎
@@
Up to equivalence, there are exactly three special pure gradings on a simple Lie algebra of type $E_{6}$ over an algebraically closed field of characteristic 00. Their universal groups and types are given in the next table:
@@
Here the quadratic four-dimensional space $\overline{E}^{\perp}/\overline{E}$ is regular and its Arf invariant is $1$, so it contains $10$ nonisotropic elements, of the form $\bar{\alpha}+\overline{E}$, with $\alpha\in\Phi^{+}$ and $\bar{\alpha}\in\overline{E}^{\perp}$. Each such left coset contains two nonisotropic vectors. The remaining $16=36-20$ positive roots give rise to one-dimensional homogeneous components and the type is $(32,20,0,0,0,1)$.
@@
In this case the quadratic two-dimensional space $\overline{E}^{\perp}/\overline{E}$ is regular and its Arf invariant is $1$, so it contains $3$ nonisotropic elements, of the form $\bar{\alpha}+\overline{E}$, with $\alpha\in\Phi^{+}$ and $\bar{\alpha}\in\overline{E}^{\perp}$. Each such left coset contains four nonisotropic vectors. The remaining $24=36-12$ positive roots $\alpha$ satisfy that the left coset $\bar{\alpha}+\overline{E}$ contains two nonisotropic vectors, and hence give rise to two-dimensional homogeneous components. The type is $(0,24,0,6,0,1)$. ∎
@@ 5. Weyl groups of the special pure gradings on simple Lie algebras of types $E_{6}$, $E_{7}$, or $E_{8}$
This section is devoted to compute the Weyl groups of the special pure gradings in the simple Lie algebras of type $E$ and to describe its action on the corresponding universal groups.
The first result provides the structure of the group $W_{s}(\Gamma)$ in Theorem 3.4.
Let $\Gamma$ be a special pure grading on a simple Lie algebra $\mathcal{L}$ of type $E_{6}$, $E_{7}$, or $E_{8}$, with universal group $G$. Let $s\in G$ be an element such that $\mathcal{L}_{s}$ contains a Cartan subalgebra $\mathcal{H}$. Let $\Phi$ be the root system relative to $\mathcal{H}$ and let $R=\mathbb{Z}\Phi$ the root space. Let $\pi\colon R\rightarrow\overline{G}$ be the projection in (3.2) and let $q\colon\overline{R}\rightarrow\mathbb{F}_{2}$ be the quadratic form in (2.2). Let $E$ be the complementary lattice of $\Gamma$ and $\overline{E}=E/2R$. Denote by $\overline{\pi}\colon\overline{R}\rightarrow\overline{G}$ the projection induced from $\pi$ and let $H$ be the subgroup of $G$ such that $\overline{H}:=H/\langle s\rangle=\overline{\pi}(\overline{E}^{\perp})$.
The quadratic form $q$ restricts to a quadratic form $q_{\overline{E}}\colon\overline{E}^{\perp}\rightarrow\mathbb{F}_{2}$. As $q$ is regular, $\overline{E}$ is the radical of $q_{\overline{E}}$. Then $q_{\overline{E}}$ is transferred by means of $\overline{\pi}$ to a regular quadratic form $q_{H,s}\colon\overline{H}\rightarrow\mathbb{F}_{2}$ (recall that $G$ is an elementary $2$-group and hence it can be seen as a vector space over $\mathbb{F}_{2}$, and the same happens with $\overline{H}$).
@@
Let $\Gamma$ be a special pure grading on a simple Lie algebra $\mathcal{L}$ of type $E_{6}$, $E_{7}$, or $E_{8}$, over an algebraically closed field of characteristic 00, with universal group $G$. With the notations preceding the theorem, we have
@@
Choose subspaces $\overline{E}^{\prime}$ and $\overline{W}$ of $\overline{R}$, such that $\overline{R}=\overline{E}\oplus\overline{E}^{\prime}\oplus\overline{W}$, where $\overline{E}^{\prime}$ is totally isotropic, $\overline{E}\cap\overline{E}^{\prime}=0$, $b_{q}|_{\overline{E}\oplus\overline{E}^{\prime}}$ is nondegenerate, and $\overline{W}$ is the orthogonal complement to $\overline{E}\oplus\overline{E}^{\prime}$ relative to $b_{q}$. To do this, it suffices to use that any isotropic vector is contained in a hyperbolic plane, and use induction on $\dim\overline{E}$ (see, e.g., the proof of [EKM08, Lemma 8.10]).
Identify $R/E\simeq\overline{R}/\overline{E}$ with $\overline{E}^{\prime}\oplus\overline{W}$, and $G$ with $(\overline{E}^{\prime}\oplus\overline{W})\times\mathbb{F}_{2}$ by means of (3.5). Then we must prove that $W_{s}(\Gamma)$, seen as a group of transformations of $(\overline{E}^{\prime}\oplus\overline{W})\times\mathbb{F}_{2}$, consists of the linear maps
@@
Note that for type $E_{7}$, where $\operatorname{\mathrm{rad}}b_{q}\neq 0$, $\operatorname{\mathrm{rad}}b_{q}\cap\overline{E}=0$ (see the proof of Theorem 4.2).
Then, with the natural identifications, and relative to the decomposition $\overline{R}=\overline{E}\oplus\overline{E}^{\prime}\oplus\overline{W}$, $\operatorname{\mathrm{Stab}}_{\mathrm{O}(q)}(\overline{E})$ consists of matrices of the form
@@
For any $f\in\mathrm{GL}(\overline{E}^{\prime})$, $g\in\mathrm{O}(q|_{\overline{W}})$, and $\alpha\in\operatorname{\mathrm{Hom}}_{\mathbb{F}_{2}}(\overline{E}^{\prime},\overline{W})$, there exists $h$, $\beta$, $\gamma$ satisfying equations (5.2), (5.3), (5.4), and hence the result follows from Theorem 3.4. ∎
@@
Let $\Gamma$ be a special pure grading on a simple Lie algebra $\mathcal{L}$ of type $E_{6}$, $E_{7}$, or $E_{8}$, over an algebraically closed field of characteristic 00, with universal group $G$. With the notations preceding Theorem 5.1, $W_{s}(\Gamma)$ is, up to isomorphism, the affine orthogonal group of the restriction of $q$ to $\overline{E}^{\perp}$.
@@
The subspaces $\overline{E}$ and $\overline{E}^{\prime}$ are conjugate under $\mathrm{O}(q)$, so we may substitute $\overline{E}$ by $\overline{E}^{\prime}$. But $(\overline{E}^{\prime})^{\perp}=\overline{E}^{\prime}\oplus\overline{W}$, and the orthogonal group of $(\overline{E}^{\prime})^{\perp}$ consists of the linear maps
$\begin{pmatrix}f&0\\ \alpha&g\end{pmatrix}$
with $f\in\mathrm{GL}(\overline{E}^{\prime})$, $g\in\mathrm{O}(q|_{\overline{W}})$, and $\alpha\in\operatorname{\mathrm{Hom}}(\overline{E}^{\prime},\overline{W})$, that is, the linear maps that appear as the upper left corner in (5.1). The result follows. ∎
@@
In case the complementary lattice $E$ is trivial: $E=2R$, which is the case in Example 2.11, one has $\overline{R}=\overline{W}$ in the proof of Theorem 5.1, and hence $W_{s}(\Gamma)$ is the orthogonal group of the degenerate quadratic form $q_{G}$ defined on $G$ (a $2$-elementary group, and hence a vector space over $\mathbb{F}_{2}$), by means of $q_{G}(g):=q\bigl{(}\upsilon(g)\bigr{)}$, with $\upsilon\colon G\rightarrow\overline{R}$ in (3.10).
@@
Let $\mathcal{L}$ be a simple Lie algebra of type $E_{7}$ or $E_{8}$ over an algebraically closed field of characteristic 00. Let $\Gamma:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}_{g}$ be the unique, up to equivalence, special pure grading with universal group $G$ and with exactly three homogeneous components: $\mathcal{L}_{s}$, $\mathcal{L}_{s^{\prime}}$ and $\mathcal{L}_{s^{\prime\prime}}$, being Cartan subalgebras. Let $W(\Gamma)$ be its Weyl group, and let $W^{\prime}(\Gamma)$ be the pointwise stabilizer in $W(\Gamma)$ of $\{s,s^{\prime},s^{\prime\prime}\}$:
@@
Let $\mathcal{L}$ be a simple Lie algebra of type $E_{7}$ or $E_{8}$ over an algebraically closed field of characteristic 00. Let $\Gamma:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}_{g}$ be the unique, up to equivalence, special pure grading with universal group $G$ and with exactly three homogeneous components, being Cartan subalgebras. Then the order of the Weyl group is given in the following table:
@@
$\begin{split}\sigma&=\Psi\left(\Bigl{(}\begin{pmatrix}0&1\\ -1&0\end{pmatrix},\cdots,\begin{pmatrix}0&1\\ -1&0\end{pmatrix}\Bigr{)}\right),\\ \tau&=\Psi\left(\Bigl{(}\begin{pmatrix}\mathbf{i}&0\\ 0&-\mathbf{i}\end{pmatrix},\cdots,\begin{pmatrix}\mathbf{i}&0\\ 0&-\mathbf{i}\end{pmatrix}\Bigr{)}\right),\end{split}$ (5.7)

### v2

@@ 4. Special pure gradings on simple Lie algebras of types $E_{6}$, $E_{7}$, or $E_{8}$
Up to equivalence, there are exactly five special pure gradings on a simple Lie algebra of type $E_{8}$ over an algebraically closed field of characteristic 0. Their universal groups and types are given in the next table:
@@
The quadratic form $q\colon\overline{R}\rightarrow\mathbb{F}_{2}$ in (2.2) is regular with trivial Arf invariant. Moreover, the nonisotropic vectors of $\overline{R}$ are precisely the $120$ vectors $\bar{\alpha}:=\alpha+2R$ for $\alpha\in\Phi^{+}$. By Proposition 3.2, the complementary lattice $E$ satisfies that $\overline{E}=E/2R$ is a totally isotropic subspace of $\overline{R}$, and hence it has dimension at most $4$. Moreover, for any dimension up to $4$, any two totally isotropic subspaces of $\overline{R}$ are conjugate under the action of $\mathrm{O}(q)$. As a consequence, we get the following possibilities, according to (3.6):
@@
On the other hand, there are 12024=96120-24=96 positive roots $\alpha\in\Phi^{+}$ with $b_{q}(\bar{\alpha},\overline{E})\neq 0$, and then the left coset $\bar{\alpha}+\overline{E}$ contains just two nonisotropic vectors, thus obtaining $48$ pairs of positive roots, which give $96$ homogeneous components in (3.6) of dimension $2$. Hence, the type is $(0,96,0,12,0,0,0,1)$.
@@
Up to equivalence, there are exactly four special pure gradings on a simple Lie algebra of type $E_{7}$ over an algebraically closed field of characteristic 0. Their universal groups and types are given in the next table:
@@
The quadratic form $q\colon\overline{R}\rightarrow\mathbb{F}_{2}$ in (2.2) is regular, but odd-dimensional. The radical of $b_{q}$ is one-dimensional: $\operatorname{\mathrm{rad}}b_{q}=\mathbb{F}_{2}a$, with $q(a)=1$. Moreover, the nonisotropic vectors of $\overline{R}$ are precisely the $63$ vectors $\bar{\alpha}:=\alpha+2R$ for $\alpha\in\Phi^{+}$ together with $a$. Any subspace of $\overline{R}$ containing strictly $\operatorname{\mathrm{rad}}b_{q}$ contains other nonisotropic vectors besides $a$, so it is not of the form $\overline{E}=E/2R$ with $E$ the complementary lattice of a special pure grading. Also, the lattice $F$ with $\overline{F}=\operatorname{\mathrm{rad}}b_{q}$ cannot be the complementary lattice of a special pure grading, graded by its universal group, because for any two different positive roots $\alpha,\beta\in\Phi^{+}$, $\bar{\alpha}-\bar{\beta}\neq a$, as otherwise we would have $1=q(\bar{\beta})=q(\bar{\alpha}+a)=q(\bar{\alpha})+q(a)=1+1=0$, and then (3.4) is not fulfilled. Therefore, the complementary lattice $E$ of a special pure grading satisfies that $\overline{E}$ is a totally isotropic subspace of $\overline{R}$ and, in particular, it does not contain $\operatorname{\mathrm{rad}}b_{q}$.
@@
Here the quadratic space $\overline{E}^{\perp}/\overline{E}$ is regular of dimension $5$, so it contains $16$ nonisotropic elements, one of them being $a+\overline{E}$. For each $\alpha\in\Phi^{+}$ with $\bar{\alpha}\in\overline{E}^{\perp}$, the coset $\bar{\alpha}+\overline{E}$ consists of two nonisotropic vectors, so either $\bar{\alpha}+\overline{E}=\{\bar{\alpha},\bar{\beta}\}$ for another positive root $\beta$, or $\bar{\alpha}+\overline{E}=\{\bar{\alpha},a\}$, which happens only for $\bar{\alpha}=a+u$, with $\overline{E}=\mathbb{F}_{2}u$. Hence there are $15$ pairs of positive roots that come together in (3.6). The other 35=633035=63-30 positive roots give rise to one-dimensional homogeneous components and the type is $(66,30,0,0,0,0,1)$.
@@
In this case the quadratic space $\overline{E}^{\perp}/\overline{E}$ is regular of dimension $3$, so it contains $4$ nonisotropic elements, one of them being $a+\overline{E}$. For each $\alpha\in\Phi^{+}$ with $\bar{\alpha}\in\overline{E}^{\perp}$, the coset $\bar{\alpha}+\overline{E}$ consists of four nonisotropic vectors, which come from four positive roots unless $\bar{\alpha}+\overline{E}=a+\overline{E}$ that contains three nonisotropic vectors obtained from positive roots. For the remaining 48=631548=63-15 positive roots $\alpha$ with $\bar{\alpha}\not\in\overline{E}^{\perp}$, the coset $\bar{\alpha}+\overline{E}$ contains two nonisotropic vectors. The type is then $(0,48,2,6,0,0,1)$.
@@
For the remaining 56=63756=63-7 positive roots $\alpha$ with $\bar{\alpha}\not\in\overline{E}^{\perp}$, the coset $\bar{\alpha}+\overline{E}$ contains four nonisotropic vectors. The type is then $(0,0,0,28,0,0,3)$. ∎
@@
Up to equivalence, there are exactly three special pure gradings on a simple Lie algebra of type $E_{6}$ over an algebraically closed field of characteristic 0. Their universal groups and types are given in the next table:
@@
Here the quadratic four-dimensional space $\overline{E}^{\perp}/\overline{E}$ is regular and its Arf invariant is $1$, so it contains $10$ nonisotropic elements, of the form $\bar{\alpha}+\overline{E}$, with $\alpha\in\Phi^{+}$ and $\bar{\alpha}\in\overline{E}^{\perp}$. Each such left coset contains two nonisotropic vectors. The remaining 16=362016=36-20 positive roots give rise to one-dimensional homogeneous components and the type is $(32,20,0,0,0,1)$.
@@
In this case the quadratic two-dimensional space $\overline{E}^{\perp}/\overline{E}$ is regular and its Arf invariant is $1$, so it contains $3$ nonisotropic elements, of the form $\bar{\alpha}+\overline{E}$, with $\alpha\in\Phi^{+}$ and $\bar{\alpha}\in\overline{E}^{\perp}$. Each such left coset contains four nonisotropic vectors. The remaining 24=361224=36-12 positive roots $\alpha$ satisfy that the left coset $\bar{\alpha}+\overline{E}$ contains two nonisotropic vectors, and hence give rise to two-dimensional homogeneous components. The type is $(0,24,0,6,0,1)$. ∎
@@ 5. Weyl groups of the special pure gradings on simple Lie algebras of types $E_{6}$, $E_{7}$, or $E_{8}$
This section is devoted to computing the Weyl group of each special pure grading on the simple Lie algebras of types $E_{6}$, $E_{7}$, $E_{8}$ and to describing its action on the corresponding universal group.
Let $\Gamma$ be a special pure grading on a simple Lie algebra $\mathcal{L}$ of type $E_{6}$, $E_{7}$, or $E_{8}$, with universal group $G$. Let $s\in G$ be an element such that $\mathcal{L}_{s}$ contains a Cartan subalgebra $\mathcal{H}$. As the first result of this section, we are going to describe the structure of the group $W_{s}(\Gamma)$ in Theorem 3.4.
Let $\Phi$ be the root system relative to $\mathcal{H}$, let $R=\mathbb{Z}\Phi$ be the root lattice, and consider the quotient $\overline{R}=R/2R$. Let $\pi\colon R\rightarrow\overline{G}$ be the projection in (3.2) and let $q\colon\overline{R}\rightarrow\mathbb{F}_{2}$ be the quadratic form in (2.2). Let $E$ be the complementary lattice of $\Gamma$ and $\overline{E}=E/2R$. Denote by $\overline{\pi}\colon\overline{R}\rightarrow\overline{G}$ the projection induced from $\pi$ (recall that $G$ is an elementary $2$-group) and let $H$ be the subgroup of $G$ such that $\overline{H}:=H/\langle s\rangle=\overline{\pi}(\overline{E}^{\perp})$.
The quadratic form $q$ restricts to a quadratic form $q_{\overline{E}}\colon\overline{E}^{\perp}\rightarrow\mathbb{F}_{2}$. As $q$ is regular and $\overline{E}$ is totally isotropic (see Section 4), $\overline{E}$ is the radical of $q_{\overline{E}}$. Hence $q_{\overline{E}}$ is transferred by means of $\overline{\pi}$ to a regular quadratic form $q_{H,s}\colon\overline{H}\rightarrow\mathbb{F}_{2}$ (where we regard $\overline{H}$ as a vector space over $\mathbb{F}_{2}$).
@@
Let $\Gamma$ be a special pure grading on a simple Lie algebra $\mathcal{L}$ of type $E_{6}$, $E_{7}$, or $E_{8}$, over an algebraically closed field of characteristic 0, with universal group $G$. With the notations preceding the theorem, we have
@@
Since $\overline{E}$ is totally isotropic, we can choose a totally isotropic subspace $\overline{E}^{\prime}$ of $\overline{R}$ such that $\overline{E}\cap\overline{E}^{\prime}=0$ and $b_{q}|_{\overline{E}\oplus\overline{E}^{\prime}}$ is nondegenerate. To do this, it suffices to use induction on $\dim\overline{E}$ and the fact that any isotropic vector is contained in a hyperbolic plane (see, e.g., the proof of [EKM08, Lemma 8.10]). Letting $\overline{W}$ be the orthogonal complement to $\overline{E}\oplus\overline{E}^{\prime}$ relative to $b_{q}$, we get $\overline{R}=\overline{E}\oplus\overline{E}^{\prime}\oplus\overline{W}$.
Identify $R/E\simeq\overline{R}/\overline{E}$ with $\overline{E}^{\prime}\oplus\overline{W}$, and $G$ with $(\overline{E}^{\prime}\oplus\overline{W})\times\mathbb{F}_{2}$ by means of (3.5). In this way, $H$ corresponds to $\overline{W}\times\mathbb{F}_{2}$. Then we must prove that $W_{s}(\Gamma)$, seen as a group of transformations of $(\overline{E}^{\prime}\oplus\overline{W})\times\mathbb{F}_{2}$, consists of the linear maps
@@
Relative to the decomposition $\overline{R}=\overline{E}\oplus\overline{E}^{\prime}\oplus\overline{W}$, the group $\operatorname{\mathrm{Stab}}_{\mathrm{O}(q)}(\overline{E})$ consists of matrices of the form
@@
For any $f\in\mathrm{GL}(\overline{E}^{\prime})$, $g\in\mathrm{O}(q|_{\overline{W}})$, and $\alpha\in\operatorname{\mathrm{Hom}}_{\mathbb{F}_{2}}(\overline{E}^{\prime},\overline{W})$, there exist $h$, $\beta$, $\gamma$ satisfying equations (5.2), (5.3), (5.4), and hence the result follows from Theorem 3.4. ∎
@@
If $\mathcal{L}$ in Theorem 5.1 has type $E_{6}$ or $E_{8}$, then $W_{s}(\Gamma)$ is isomorphic to the affine orthogonal group of the restriction of $q$ to $\overline{E}^{\perp}$.
@@
Note that $\overline{E}^{\perp}=\overline{E}\oplus\overline{W}$. Moreover, for $E_{6}$ and $E_{8}$, $\operatorname{\mathrm{rad}}b_{q}$ is trivial, so the orthogonal group $\mathrm{O}(q|_{\overline{E}^{\perp}})$ is isomorphic to the image of $\bar{\rho}$ by means of the map
[…]
for $h\in\mathrm{GL}(\overline{E})$, $g\in\mathrm{O}(q|_{\overline{W}})$ and $\gamma\in\operatorname{\mathrm{Hom}}(\overline{W},\overline{E})$, where $f\in\mathrm{GL}(\overline{E}^{\prime})$ is determined by (5.2) and $\alpha$ by (5.3).
On the other hand, $\overline{E}^{\perp}$ is naturally isomorphic to $\operatorname{\mathrm{Hom}}(\overline{R}/\overline{E},\mathbb{F}_{2})$ by means of $b_{q}$. Denote this isomorphism by $\theta$.
As we have seen, Theorem 3.4 shows that $W_{s}(\Gamma)$ is isomorphic to the semidirect product $\operatorname{\mathrm{Hom}}(R/E,\mathbb{F}_{2})\rtimes\operatorname{\mathrm{im}\,}\bar{\rho}\leq\operatorname{\mathrm{Hom}}(R/E,\mathbb{F}_{2})\rtimes\mathrm{GL}(R/E)$, where the action of $\mathrm{GL}(R/E)$ on $\operatorname{\mathrm{Hom}}(R/E,\mathbb{F}_{2})$ is the natural one: $f.\xi=\xi\circ f^{-1}$. We claim that our isomorphisms $\Theta$ and $\theta$ are compatible with the action, i.e., $\theta(Ax)=\Theta(A).\theta(x)$ for all $A\in\mathrm{O}(q|_{\overline{E}^{\perp}})$ and $x\in\overline{E}^{\perp}$. Using the identification $R/E\simeq\overline{E}^{\prime}\oplus\overline{W}$, this is equivalent to the equation $b_{q}(A(u+w),\Theta(A)(v^{\prime}+w^{\prime}))=b_{q}(u+w,v^{\prime}+w^{\prime})$ for all $u\in\overline{E}$, $v^{\prime}\in\overline{E}^{\prime}$, $w,w^{\prime}\in\overline{W}$, and $A=\left(\begin{smallmatrix}h&\gamma\\ 0&g\end{smallmatrix}\right)\in\mathrm{O}(q|_{\overline{E}^{\perp}})$, which is easy to check:
$\begin{split}b_{q}\bigl(A(u+w),\Theta(A)(v^{\prime}+w^{\prime})\bigr)&=b_{q}\bigl(h(u)+\gamma(w)+g(w),f(v^{\prime})+\alpha(v^{\prime})+g(w^{\prime})\bigr)\\ &=b_{q}\bigl(h(u)+\gamma(w),f(v^{\prime})\bigr)+b_{q}\bigl(g(w),\alpha(v^{\prime})+g(w^{\prime})\bigr)\\ &=b_{q}(u,v^{\prime})+b_{q}(w,w^{\prime})=b_{q}(u+w,v^{\prime}+w^{\prime}),\end{split}$
where we have used (5.2) and (5.3).
We conclude that $W_{s}(\Gamma)$ is isomorphic to the semidirect product $\overline{E}^{\perp}\rtimes\mathrm{O}(q|_{\overline{E}^{\perp}})$, which is the affine orthogonal group of the restriction of $q$ to $\overline{E}^{\perp}$. ∎
@@
In case the complementary lattice $E$ is trivial: $E=2R$, which is the case in Example 2.11, one has $\overline{R}=\overline{W}$ in the proof of Theorem 5.1, and hence $W_{s}(\Gamma)$ is the orthogonal group of the degenerate quadratic form $q_{G}$ defined on $G$ (a 22-elementary group, and hence a vector space over $\mathbb{F}_{2}$), by means of $q_{G}(g):=q\bigl(\upsilon(g)\bigr)$, with $\upsilon\colon G\rightarrow\overline{R}$ in (3.10).
@@
Let $\mathcal{L}$ be a simple Lie algebra of type $E_{7}$ or $E_{8}$ over an algebraically closed field of characteristic 0. Let $\Gamma:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}_{g}$ be the unique, up to equivalence, special pure grading with universal group $G$ and with exactly three homogeneous components: $\mathcal{L}_{s}$, $\mathcal{L}_{s^{\prime}}$ and $\mathcal{L}_{s^{\prime\prime}}$, being Cartan subalgebras. Let $W(\Gamma)$ be its Weyl group, and let $W^{\prime}(\Gamma)$ be the pointwise stabilizer in $W(\Gamma)$ of $\{s,s^{\prime},s^{\prime\prime}\}$:
@@
Let $\mathcal{L}$ be a simple Lie algebra of type $E_{7}$ or $E_{8}$ over an algebraically closed field of characteristic 0. Let $\Gamma:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}_{g}$ be the unique, up to equivalence, special pure grading with universal group $G$ and with exactly three homogeneous components being Cartan subalgebras. Then the order of the Weyl group is given in the following table:
@@
[…]